Q) a gun kept at a straight horz. road is used to hit a car , travelling along the same road away from the gun with a uniform speed 72 x 21/2 km/hr . the car is at a distance of 50 m from the gun , where the gun is fired at an angle of 45o with the horizontal. find 

1)the distance of car from the gun when the shell hits it.

2)the speed of projection of the shell from the gun. 

Grade:11

1 Answers

Arun
6 years ago
Speed = 722 × 100060 × 60 = 722 × 518 = 202 m s- 1. The distance moved by the car in time t = 202 × t. The distance moved by the shell of gun in time t = u cos 45 × t .the time when shell hits the car, the distance moved is given by, 50 + 202 t = u × 12 × t. - - - 1t = 2u sin 45g = 2ug2 - - - 2 from eqn 1 and 2, we get, 50 + 202 × 2 ug2 = u × 12 × 2ug 2 or 50 + 4u = u 210 = u2 - 40u - 500. u = 40 ± 1600 + 20002 or u = 20 ± 30. or u = 50m s - 1. therefore distance when shell hits the car= u210 = 250010 = 250m.